OpenStudy (anonymous):

B=h+w ----- Solve for h 2

OpenStudy (katrinakaif):

2B - w = h As you have to multiply by the recripocal of 1/2 which is 2/1 onto the other side.. so 2/1 x B = h + w Then substract w from both sides to get 2b - w = h